OpenStudy (anonymous):

the formula A=p(1+r)^2 gives the amount A in dollars that P dollars will grow to in 2 yrs. at interest rate r using compound interest. what interest rate will cause $2,000 to grow to $2,163.20 in 2 yrs.

OpenStudy (turingtest):

A=2163.2 p=2000 so we need to solve for the remaining variable, r: A=p(1+r)^2 (1+r)^2=A/p \[r=\sqrt{A/p}-1=\sqrt{2163.2/2000}-1=0.04\] or a 4% interest rate.